High temperatures in a certain city for the month of August follow a uniform distribution over the interval 66°F to 88°F. What is the probability that a randomly selected August day has a high temperature that exceeded 71°F? (Round your answer to 4 decimal places)
Answer by greenestamps(13198)



The temperatures are evenly distributed over the continuous interval from 66 to 88 degrees.

The "good" temperatures are uniformly distributed over the continuous interval from 71 to 88 degrees.

P(temperature above 71 degrees) = (88-71)/(88-66) = 17/22.

You can do the conversion to a decimal answer....